OpenStudy (anonymous):

using partial fractions. x/x-1. x^2/x-1.

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

the question doesn't make sense :\

OpenStudy (unklerhaukus):

is this your question?\[y(x)=\frac{x}{x-1}\qquad\qquad\qquad w(x)=\frac{x^2}{x-1}\] \[\frac{\partial y}{\partial x}=?\qquad\qquad\qquad \frac{\partial w}{\partial x}=?\]

OpenStudy (unklerhaukus):

oh, well these are partial derivatives , partial fractions is different thing

OpenStudy (anonymous):

it's not like y is a function of more than 1 variable, is it? Why the 'partial' derivative?

OpenStudy (anonymous):

sorry man I get confused, actually the teacher told us to solve these questions .. & we're studying partial fractions.

OpenStudy (unklerhaukus):

ignore what i wrote earlier then

OpenStudy (anonymous):

so any help*?

OpenStudy (unklerhaukus):

i dunno